# Automated Detection of Galactic Rings from SDSS Images

### v1

@@ 8 A Catalog of Ringed galaxies with Bars
Bars in galaxies are much more common than rings. Approximately half of the observed galaxies are seen to have bars Knapen (1999). However galaxies that contain both bars and rings are certainly a smaller fraction. Dedicated catalogs for such galaxies are also quite rare. Here we use the Abraham et al. (2018) catalog, which is an automated catalog of galaxies with bars in SDSS, to find galaxies with both bars and rings. The Abraham et al. (2018) catalog contains 111,838 galaxies that have been predicted as either barred or unbarred. We selected the barred galaxies which came to be 25,781 in number and used our trained network to predict which of these galaxies also have rings. Figure 13 shows a random selection of the barred galaxies that have been predicted to have rings with a probability score of greater than 90 percent. We found 2087 galaxies that contained both bars and rings using a classification threshold of $0.5$. A preview of the catalog containing barred galaxies with predictions for the presence or absence of a ring is shown in Table 1.
Refer to caption Galaxies with rings
Refer to caption Galaxies with bars and rings
Figure 13: Some good examples of ringed galaxies (above) and barred ring galaxies(below) from our catalog.
@@ 9 Conclusion
We have shown that a relatively simple deep neural network based on the AlexNet architecture can be used to detect galactic rings from colour composite galaxy images. The training required only a modest sample size of 1122 original ringed galaxy images and about 10639 original images of galaxies without rings. To overcome the challenges of training with a small dataset of labelled rings, we used a smaller network and also increased the total number of images available for training through realistic image transformations, and during testing we used k-fold cross validation to make the best use of the data. We also used evaluation metrics suited for the class imbalance that was present in our data. We have prepared a catalog of candidate ring and non-ring galaxies using the predictions of our trained network. The probability score provided in the catalog can be used as a confidence measure of the presence of rings, to obtain smaller catalogues with the required level of confidence for follow up analysis. Additionally we used the network to generate a catalog of galaxies with rings as well a bars, which may be of special interest to galaxy morphologists.

### v2

@@ 8 A Catalog of Ringed galaxies with Bars
Bars in galaxies are much more common than rings. Approximately half of the observed galaxies are seen to have bars (Knapen, 1999). However galaxies that contain both bars and rings are certainly a smaller fraction. Dedicated catalogs for such galaxies are also quite rare. Here we use the Abraham et al. (2018) catalog, which is an automated catalog of galaxies with bars in SDSS, to find galaxies with both bars and rings. The Abraham et al. (2018) catalog contains 111,838 galaxies that have been predicted as either barred or unbarred. We selected the barred galaxies which came to be 25,781 in number and used our trained network to predict which of these galaxies also have rings. Figure 10 shows a random selection of the barred galaxies that have been predicted to have rings with a probability score of greater than 90 percent. We found 2087 galaxies that contained both bars and rings using a classification threshold of $0.5$. A preview of the catalog containing barred galaxies with predictions for the presence or absence of a ring is shown in Table 1.
9 Discussion
As a demonstration of the usefulness of the catalogs generated in this paper, in this section, we will explore the environments and star formation characteristics of ring galaxies in our sample. Our goal is to analyze star formation within the framework of the main sequence of star-forming galaxies, the transitioning population of green valley galaxies, and the quenched galaxies. Additionally, we will examine star formation in relation to the environmental variations from low to high densities. For this analysis, we use a sample of 4855 galaxies with a classification threshold of 0.90 that is predicted to have rings.
[…]
Figure 14 presents the histogram of $\log\ \Sigma$ for the sample of 4471 ring galaxies with environment information. We find that ring galaxies are primarily found in intermediate density environments, i.e. in galaxy groups (49.3%) and in low-density environments, i.e. as isolated galaxies 35.1%). We see relatively few ring galaxies (15.6%) in high-density environments in agreement with Elmegreen et al. (1992). Numerical simulations, such as ROMULUS (Tremmel et al., 2017), suggest that galaxies within group environments often undergo interactions or multiple mergers with gas-rich satellites, evolving into elliptical galaxies. Subsequently, these galaxies start to accrete gas and redevelop their disks, which then become unstable, triggering star formation in ring-like patterns. When galaxies display rejuvenated star-forming rings, two pathways for gas inflow are identified in numerical simulations: (a) Diffuse cooling gas that feeds the galaxy, aiding in developing the disk/ring structure. (b)The acquisition of ram pressure-stripped gas streams from gas-rich satellites (Jung et al., 2022).
To study star formation in our sample ring galaxies with environment information, we utilized the GALEX-SDSS-WISE Legacy Catalog (GSWLC-X2) (Salim et al., 2016, 2018) and the cross-match gave us 4022 ring galaxies with a stellar mass (log M) range of 8.8-11.68. To investigate the influence of the environment on the star formation process in ring galaxies, we constructed a control sample of non-ring galaxies with a classification threshold of 0.40 for being a ring galaxy, i.e., a threshold of 0.60 for being a non-ring galaxy. This control sample has a similar r-band magnitude and redshift, and a cross-match with the GSWLC-X2 catalog gave $\sim$5000 non-ring galaxies. The GSWLC-X2 catalog integrates ultraviolet data from GALEX, optical data from SDSS, and infrared data from the WISE all-sky survey to model galaxy spectral energy distribution (SED). During SED modelling, the catalog applies flexible dust attenuation laws and emission line corrections to ensure accurate measurements of star formation rate (SFR) and stellar mass (M), as it might lead to the wrong identification of galaxies.
We calculated the specific star formation rate (sSFR) and, following Salim (2014), classified our sample galaxies into three sub-classes as a function of environment, as shown in Figure 15. The star-forming region is identified with log sSFR $\geq$ -10.8, the green valley, which is a transition zone between the star-forming and quenched regions, is defined as -11.8 < log sSFR < -10.8, and quenched region is defined by log sSFR $\leq$ -11.8. In high-density environments, most of the ring galaxies are in the quenched (42.3 %) or green valley region (33.6 %); however, the control sample of non-ring galaxies are almost equally distributed in quenched (33.8%) and star-forming region (47.8%) with fewer galaxies in green valley (18.9%). In contrast, low and intermediate-density environments are dominated by star-forming ring galaxies (52.4% and 43.2 %, respectively), as the non-ring galaxies from the control sample (86.9% and 76.3 %, respectively). However, the number of ring galaxies in green valleys increases for intermediate-density environments, though the number of non-ring galaxies increases for high-density environments. It is apparent that most of the ring galaxies are located in the star-forming region for low and intermediate-density environments, which validates the scenario in which rejuvenation of star formation happens by accreting gas and triggering star formation in ring-like features. Our results are somewhat different from those of Fernandez et al. (2024), who suggest that ringed galaxies exhibit lower star formation activity in intermediate-density environments, though it is much lower in high-density environments.
Many ring galaxies from our sample are also part of the MaNGA survey, and the study of spatially resolved stellar populations of rings using such IFU surveys will provide useful insight into the origin of star-forming rings, and this can be the subject of future work.
Figure 14: Histogram of local environmental density ($\log\ \Sigma\ (\text{Mpc}^{-2})$) for our final sample of galaxies. The dashed lines separate the sample into low ($\log\Sigma\ (\text{Mpc}^{-2})<-0.5$), intermediate(-0.5 < $\log\Sigma\ (\text{Mpc}^{-2})<0.5$), and high densities($\log\Sigma\ (\text{Mpc}^{-2})>0.5$).
Figure 15: Dependence of specific star formation rate on stellar mass (sSFR-Mass plane) for our sample of ring galaxies and control sample of non-ring galaxies (gray dot) as a function of (a) low-density ($\log\Sigma\ (\text{Mpc}^{-2})<-0.5$) (b) intermediate-density (-0.5 < $\log\Sigma\ (\text{Mpc}^{-2})<0.5$), and (c) high-density environment ($\log\Sigma\ (\text{Mpc}^{-2})>0.5$)
@@ 10 Conclusion
We have shown that a relatively simple deep neural network based on the AlexNet architecture can be used to detect galactic rings from colour composite galaxy images. The training required only a modest sample size of 1122 original ringed galaxy images and about 10639 original images of galaxies without rings. To overcome the challenges of training with a small dataset of labelled rings, we used a smaller network and also increased the total number of images available for training through realistic image transformations. We also used evaluation metrics suited for the class imbalance that was present in our data. We have prepared a catalog of candidate ring and non-ring galaxies using the predictions of our trained network. The probability score provided in the catalog can be used as a confidence measure of the presence of rings, to obtain smaller catalogs with the required level of confidence for follow up analysis. Additionally we used the network to generate a catalog of galaxies with rings as well a bars, which may be of special interest to galaxy morphologists. Finally, we explored the connection between the environments and star formation characteristics of the 4855 ringed galaxies in our catalog. We found that ring galaxies are star-forming and primarily found in low- and intermediate-density environments.